OpenStudy (anonymous):

solve the system of equations below by graphing them with a pencil and paper. enter your answer as an ordered pair. y=-3x+1. y=-x+5

OpenStudy (anonymous):

@AnswerMyQuestions help me plzz

OpenStudy (anonymous):

sorry,i was wrong ,i will rewrite

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so i make it up

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

@song6

OpenStudy (anonymous):

X=-2,Y=7

OpenStudy (anonymous):

@song6 I need help with a lot of them

OpenStudy (anonymous):

this one is the same question but different numbas

OpenStudy (anonymous):

y=-x+5 y=x-3

OpenStudy (anonymous):

X=4.Y=1